# The Core of Approval-Based Committee Elections with Few Candidates

### v1

@@ 1 Introduction
It might seem surprising that the state of the art hasn’t improved further than that. In particular, there is a natural way of using mixed integer linear program to search for counterexamples to core existence: fix $m$ and $k$, and introduce a fractional variable for each possible ballot, indicating what fraction of the voters submit this ballot. Then, for every possible committee, enforce using binary variables that there exists at least one successful core deviation. If an ILP solver determines that the resulting program is infeasible, this implies the non-emptiness of the core for $m$ and $k$, for any number $n$ of voters. Unfortunately, the size of this program grows rapidly, and these programs are not easy to solve even for very small sizes (Gurobi solves $m=7$, $k=5$, in 450s, but did not solve $m=7$, $k=4$ after 134 000s on 8 cores). [^1] The same problem can also be encoded as an SMT problem on linear arithmetic. This can sometimes lead to faster solve times, but it is also only feasible for very small sizes.
@@
These results are obtained by analyzing variants of Proportional Approval Voting (PAV), the voting rule proposed by Thiele (1895). This voting rule works by maximizing a carefully chosen objective function over the set of all committees of size $k$. We show via linear programs that PAV always selects a core-stable committee when $k\leqslant 7$, and that it always selects at least one core-stable committee when $k=8$, perhaps tied with other committees that fail core-stability. For the results about limited numbers of candidates, we consider a recursive version of PAV where, if the PAV committee fails the core because some set $T$ has too much support, we then re-compute PAV subject to the constraint that $T\subseteq W$ and without taking into account the voters who were part of the blocking coalition. If the result still fails core-stability, we add additional constraints. We show that if $m\leqslant 15$, this process always terminates with a core-stable committee.
@@ 2 Related Work
It is NP-hard to compute the PAV rule (Aziz et al., 2015, Corollary 1), but it is fixed-parameter tractable for a variety of parameters (Yang and Wang, 2023). A local search variant of PAV retains it proportionality properties and can be computed in polynomial time (Aziz et al., 2018) for an appropriately chosen tolerance parameter (Kraiczy and Elkind, 2024). The problem of checking whether a given committee is in the core is coNP-complete (Brill et al., 2022, Theorem 5.3) and remains hard even when every voter approves at most 6 candidates, and each candidate is approved by at most 2 voters (Munagala et al., 2022a, Theorem 1). The verification problem is also hard to approximate to within a factor better than $1+1/e$ (Munagala et al., 2022a, Theorem 2), though a logarithmic approximation algorithm exists (Munagala et al., 2022a, Theorem 3).
@@
Peters and Skowron (2020) show that PAV is in the $(2,0)$-core, and that the Method of Equal Shares is in the $(\log k,0)$-core (or a mild relaxation of that concept). Munagala et al. (2022b) show that the $(9.27,0)$-core is non-empty even for general additive valuations. Fain et al. (2018, Appendix C) show that the $(1+\varepsilon,0)$-core is non-empty if it is additionally additively relaxed. Jiang et al. (2020) show that the $(0,16)$-core is non-empty, via rounding a stable lottery (Cheng et al., 2020); they conjecture that at least the $(0,2)$-core is always non-empty. A similar rounding technique has been used to analyze Condorcet winning sets (Charikar et al., 2024).
@@
Analogs of core-stability have also been defined in fractional models. For example, Aziz et al. (2020) consider “fair mixing” in an approval-based model, where the output is a probability distribution over candidates (which can be interpreted as a division of a budget). They show that the rule maximizing Nash welfare (which is related to the PAV rule) satisfies core-stability. Fain et al. (2016) obtain the same result in a more general model with additive linear utilities. They also show that a core-stable outcome exists for fractional committees (which can be viewed as a probability distribution where each candidate receives mass at most $1/k$) via Lindahl equilibrium which is known to exist from fixed-point theorems (Foley, 1970). Munagala et al. (2022b) expand on the use of Lindahl equilibrium. Various weakenings of the core have also been studied in approval-based fair mixing and related models (see, e.g., Brandl et al., 2021; Suzuki and Vollen, 2024; Bei et al., 2024).
@@ 3 Preliminaries
A set $T\subseteq C$ with $|T|\leqslant k$ is called a potential deviation. Given a profile $P$, a committee $W$ is core stable if for every potential deviation $T$, we have
@@
The PAV rule selects the committees with highest PAV score.
We will be interested in local maxima of the PAV objective, i.e., committees that have a higher PAV score than any committee obtained by performing a single swap. For a committee $W$, $x\in W$ and $y\in C\setminus W$, we write $W_{xy}=W\setminus\{x\}\cup\{y\}$ for the committee obtained by replacing $x$ with $y$. Given a profile $P$ and a fixed committee $W$, we write
@@
for the increase in PAV score resulting from this swap. For a ballot $A$, we define $\Delta_{A,x,y}$ analogously. Then we say that $W$ is a PAV committee if $\Delta_{P,x,y}\leqslant 0$ for all $x$ and $y$. (We could also define “PAV committee” to refer only to committees with the highest PAV score. All results would continue to hold, but the local swap version leads to smaller linear programs.)
The following lemma shows that a PAV committee never admits deviations of certain kinds: neither disjoint deviations, nor deviations that contain only a single unelected candidate.
@@
Let $W$ be a PAV committee, and let $T$ be a potential deviation. Suppose we have
@@
(ii) If $T\setminus W=\emptyset$, then $T\subseteq W$, and there are no approval ballots that strictly prefer $T$ to $W$. If $T\setminus W=\{c\}$, then for every $A\in\mathcal{A}$ with $u_{A}(T)>u_{A}(W)$, we have $c\in A$. Writing $\ell=|T|$, if $T$ were a deviation, we would thus have a set $S$ of ballots forming $\ell/k$ of the profile, who all approve $c\not\in W$, and who all have utility $u_{A}(W)<u_{A}(T)\leqslant\ell$. Thus, we have a violation of the EJR+ axiom of Brill and Peters (2023) which PAV committees are known to satisfy. ∎
@@ 4.1 Committee size $k\leqslant 7$
For committee sizes up to $7$, core-stable committees always exist because every PAV committee is in the core. This establishes existence of core-stable committees, but also indicates that PAV rule is a very good rule for smaller committee sizes. [^2] There are examples where the sequential Phragmén rule fails core (and even EJR) for $k=6$, and where MES fails core for $k=7$. These counterexamples work even for the party-approval setting (Brill et al., 2022), where each candidate can be placed in the committee several times. For Phragmén, take the 3-voter profile $(ab,bc,ac)$, where Phragmén can elect $ababab$, with $T=\{c,c,c,c\}$ forming a deviation. For MES, take the 7-voter profile $(ab,ac,ad,bcd,bcd,bcd,bcd)$, where $bbbbbaa$ is an outcome of MES, with $T=\{c,c,c,d,d,d\}$ forming a deviation. (In these examples, there are other tied outcomes in the core, and I don’t know if unique examples exist.) Like other proofs that PAV is proportional, our proof reasons about the change in PAV score caused by certain swaps. To establish a key inequality, the proof refers to the result of a computer enumeration of all possible ballot types, which can be reproduced using Python code available on GitHub.
@@
When $k\leqslant 7$, every PAV committee is in the core.
@@
We now consider the following process: we remove some candidate $x\in W\setminus T$ from the committee $W$, and replace it by some candidate $y\in T\setminus W$, and will estimate the total effect of these swaps. Because $W$ is a PAV committee, we have
@@
$a=(W\setminus T)\cap A,\quad b=(W\cap T)\cap A,\quad c=(T\setminus W)\cap A.$
@@
While it may not look like it, the result of Theorem 4.1 was obtained by linear programming. Suppose that the result is false, so that there exists a profile $P$ and a PAV committee $W$ such that some $T$ is a successful deviation. Without loss of generality, there exists such an example with $W=\{c_{1},\dots,c_{k}\}$. Note that $P$ then forms a solution to the following system of linear inequalities, where we may assume that $C=W\cup T$: [^4] We may make this assumption because a counterexample on a larger $C$ remains a counterexample when restricted to $W\cup T$, because a PAV committee remains a PAV committee after deleting candidates outside the committee.
@@
Due to the following remark, it is possible to compute a core-stable outcome in polynomial time whenever $k\leqslant 7$.
@@
Let $k\leqslant 7$ and $\varepsilon=0.1/k^{2}$. By solving linear programs [GitHub], one can check that every committee that is $\varepsilon$-local-swap-stable (i.e., $\Delta_{P,x,y}\leqslant\varepsilon$ for all $x$ and $y$) is in the core. An $\varepsilon$-local-swap-stable and thus a core-stable committee can be found in polynomial time (see Aziz et al., 2018, Proposition 1). Note that the statement is not true for $\varepsilon=1/k^{2}$, even though this $\varepsilon$ is enough to ensure that the committee satisfies EJR (Aziz et al., 2018, Theorem 1). For example, in the profile with $P(\{a,b\})=P(\{a,c\})=0.25$ and $P(\{d,e,f,g,h\})=0.5$, for $k=6$, the committee $\{a,d,e,f,g,h\}$ fails the core due to $T=\{a,b,c\}$, but it is $1/40$-local-swap-stable, and $1/40<1/36=1/k^{2}$.
@@ 4.2 Committee size $k=8$
Theorem 4.1 does not hold for $k=8$: There are profiles where some PAV committee is not in the core.
@@
On this profile, $W=\{c_{1},c_{2},c_{5},c_{6},c_{7},c_{8},c_{9},c_{10}\}$ is a PAV committee (indicated in blue in the picture). However, $W$ is not in the core: consider $T=\{c_{1},c_{2},c_{3},c_{4}\}$, which has support from $\frac{1}{2}$ of the voters, and $|T|/k=\frac{1}{2}$. ∎
Note, however, that in Example 4.3, there is more than one PAV committee. In particular, $W^{\prime}=\{c_{1},\dots,c_{8}\}$ is also a PAV committee and it is in the core. (Both of these committees are selected by the PAV rule.)
@@
Let $P$ be a profile and suppose that $W$ with $|W|=8$ is a PAV committee that is not in the core due to objection $T$. Then there exist distinct $a,b\in W$ and distinct $x,y\in C\setminus W$ such that $T=\{a,b,x,y\}$. In addition,
@@
We first check that if $W$ is not in the core, then any core objection must use a $T$ with $|T|=4$ and $|W\cap T|=2$. This can be deduced using the linear programming approach behind Theorem 4.1; by iterating through all possible $T$, we find that the system (4) has a solution only for $T$ satisfying the condition in the theorem statement. Alternatively, one can check [GitHub] that only for such $T$ can the inequality (3) be violated; thus for other $T$ the proof of Theorem 4.1 goes through.
Now fix such a $T=\{a,b,x,y\}$. Assume that there exists a profile $P$ that violates where $W$ is a PAV committee with successful deviation $T$ but that violates any of the conditions (i)–(iii). Then if we delete all candidates outside $W\cup T$ from the profile, it would still fail (i)–(iii). Thus, for purposes of making the following linear programs finite, we may assume that $C=W\cup T$ (so $|C|=10$).
@@
As we discussed, Example 4.3 shows an example of a PAV committee that is not core-stable, but there are other PAV committees for the same profile that are core-stable. Thanks to Lemma 4.4, we deduce that the same holds for all counterexamples. Hence, for every instance, at least one PAV committee is in the core, and thus the core is always non-empty for $k=8$.
@@
When $k=8$, some PAV committee is in the core.
@@
Let $k=8$ and let $P$ be a profile. If on $P$, all PAV committees are in the core, we are done. So suppose that $W$ is a PAV committee that is not core stable due to objection $T$. From Lemma 4.4, there exist distinct $a,b\in W$ and distinct $x,y\in C\setminus W$ such that $T=\{a,b,x,y\}$. Take any $c\in W\setminus\{a,b\}$. Then the committee $W^{\prime}=W\setminus\{c\}\cup\{x\}$ has the same PAV score as $W$, because the removal of $c$ causes a decrease in PAV score of $1/12$ and the addition of $x$ causes an increase of at least $\frac{1}{4}\cdot\frac{1}{3}=1/12$ due to the quarter of voters from (i) with ballots $A$ such that $A\cap(W\cup T)=\{a,b,x\}$. Thus, $W^{\prime}$ is also a PAV committee. We now show that $W^{\prime}$ is in the core.

### v2

# The Core of Approval-Based Committee Elections with Few Seats

@@ 1 Introduction
It might seem surprising that the state of the art has not improved beyond these very small parameter values. In particular, there is a natural way of using mixed integer linear programming (ILPs) to search for counterexamples to core existence: fix $m$ and $k$, and introduce a fractional variable for each possible ballot, indicating what fraction of the voters submit this ballot. Then, for every possible committee, enforce using binary variables that there exists at least one successful core deviation. If an ILP solver determines that the resulting program is infeasible, this implies the non-emptiness of the core for $m$ and $k$, for any number $n$ of voters. Unfortunately, the size of this program grows rapidly, and they are not easy to solve even for very small sizes (Gurobi solves $m=7$, $k=5$, in 450s, but did not solve $m=7$, $k=4$ after 134 000s (37h) on 8 cores). [^1] The same problem can also be encoded as an SMT problem on linear arithmetic. This can sometimes lead to faster solve times, but it is also only feasible for very small sizes.
@@
These results are obtained by analyzing variants of Proportional Approval Voting (PAV), the voting rule proposed by Thiele (1895). This voting rule works by maximizing a carefully chosen objective function over the set of all committees of size $k$. We show via linear programs that PAV always selects a core-stable committee when $k\leqslant 7$, and that it always selects at least one core-stable committee when $k=8$, perhaps tied with other committees that fail core-stability. This performance of the PAV rule is remarkable, and contrasts with other rules such as the Phragmén rule and the Method of Equal Shares which fail the core for $k=6$ and $k=7$, respectively (see Footnote 3). In many applications, the number of seats to fill will be 8 or fewer, and thus our result suggests that PAV is a good rule for such contexts. [^2] Note, however, that even for $k=6$, PAV may select committees that fail laminar proportionality (Peters and Skowron, 2020) and are intuitively unfair. A three-voter example for $k=6$ is $A_{1}=\{c_{1},c_{2},c_{3}\}$, $A_{2}=\{c_{1},c_{2},c_{4}\}$, and $A_{3}=\{c_{5},c_{6},c_{7}\}$ where $W=\{c_{1},c_{2},c_{3},c_{5},c_{6},c_{7}\}$ is a global PAV committee (tied with others) while laminar proportionality demands that $\{c_{1},c_{2},c_{3},c_{4}\}\subseteq W$.
For the results about limited numbers of candidates, we consider a recursive version of PAV where, if the PAV committee fails the core because some set $T$ has too much support, we then re-compute PAV subject to the constraint that $T\subseteq W$ and without taking into account the voters who were part of the blocking coalition. If the result still fails core-stability, we add additional constraints. We show that if $m\leqslant 15$, this process always terminates with a core-stable committee. However, there are examples where the rule fails to guarantee the core for $m=16$.
@@ 2 Related Work
It is NP-hard to compute the PAV rule (Aziz et al., 2015, Corollary 1), but it is fixed-parameter tractable for a variety of parameters (Yang and Wang, 2023). A local search variant of PAV retains its proportionality properties and can be computed in polynomial time (Aziz et al., 2018) for an appropriately chosen tolerance parameter (Kraiczy and Elkind, 2024). The problem of checking whether a given committee is in the core is coNP-complete (Brill et al., 2022, Theorem 5.3) and remains hard even when every voter approves at most 6 candidates, and each candidate is approved by at most 2 voters (Munagala et al., 2022a, Theorem 1). The verification problem is also hard to approximate to within a factor better than $1+1/e$ (Munagala et al., 2022a, Theorem 2), though a logarithmic approximation algorithm exists (Munagala et al., 2022a, Theorem 3).
@@
Peters and Skowron (2020) show that PAV is in the $(2,1)$-core, and that the Method of Equal Shares is in the $(\log k,1)$-core (or a mild relaxation of that concept). Munagala et al. (2022b) show that the $(9.27,1)$-core is non-empty even for general additive valuations. Fain et al. (2018, Appendix C) show that the $(1+\varepsilon,1)$-core is non-empty if it is additionally additively relaxed. Jiang et al. (2020) show that the $(1,16)$-core is non-empty, via rounding a stable lottery (Cheng et al., 2020); they conjecture that at least the $(1,2)$-core is always non-empty. A similar rounding technique has been used to analyze Condorcet winning sets (Charikar et al., 2024).
@@
Analogs of core-stability have also been defined in fractional models. For example, Aziz et al. (2020) consider “fair mixing” in an approval-based model, where the output is a probability distribution over candidates (which can be interpreted as a division of a budget). They show that the rule maximizing Nash welfare (which is related to the PAV rule) satisfies core-stability. Fain et al. (2016) obtain the same result in a more general model with additive linear utilities. They also show that a core-stable outcome exists for fractional committees (which can be viewed as a probability distribution where each candidate receives mass at most $1/k$) via Lindahl equilibrium which is known to exist from fixed-point theorems (Foley, 1970; Munagala et al., 2022b) and via convex programming (Kroer and Peters, 2025). Various weakenings of the core have also been studied in approval-based fair mixing and related models (see, e.g., Brandl et al., 2021; Suzuki and Vollen, 2024; Bei et al., 2025).
@@ 3 Preliminaries
A nonempty set $T\subseteq C$ with $|T|\leqslant k$ is called a potential deviation. Given a profile $P$, a committee $W$ is core stable if for every potential deviation $T$, we have
@@
A committee is called a global PAV committee if it has the highest PAV score. The PAV rule selects all global PAV committees.
We will also be interested in local maxima of the PAV objective, i.e., committees that have a weakly higher PAV score than any committee obtained by performing a single swap. For a committee $W$, $x\in W$ and $y\in C\setminus W$, we write $W_{xy}=W\setminus\{x\}\cup\{y\}$ for the committee obtained by replacing $x$ with $y$. Given a profile $P$ and a fixed committee $W$, we write
@@
for the increase in PAV score resulting from this swap. For a ballot $A$, we define $\Delta_{A,x,y}$ analogously. Then we say that $W$ is a local PAV committee if $\Delta_{P,x,y}\leqslant 0$ for all $x$ and $y$. Note that every global PAV committee is also a local PAV committee.
The following lemma shows that a local PAV committee never admits deviations of certain kinds: neither disjoint deviations, nor deviations that contain only a single unelected candidate.
@@
Let $W$ be a local PAV committee, and let $T$ be a potential deviation. Suppose we have
@@
(ii) If $T\setminus W=\emptyset$, then $T\subseteq W$, and there are no approval ballots that strictly prefer $T$ to $W$. If $T\setminus W=\{c\}$, then for every $A\in\mathcal{A}$ with $u_{A}(T)>u_{A}(W)$, we have $c\in A$. Writing $\ell=|T|$, if $T$ were a deviation, we would thus have a set $S$ of ballots forming $\ell/k$ of the profile, who all approve $c\not\in W$, and who all have utility $u_{A}(W)<u_{A}(T)\leqslant\ell$. Thus, we have a violation of the EJR+ axiom of Brill and Peters (2023) which local PAV committees are known to satisfy. ∎
@@ 4.1 Committee size $k\leqslant 7$
For committee sizes up to $7$, core-stable committees always exist because every local PAV committee is in the core. This establishes existence of core-stable committees, but also indicates that PAV rule is a very good rule for smaller committee sizes. [^3] There are examples where the sequential Phragmén rule fails core (and even EJR) for $k=6$, and where MES fails core for $k=7$. These counterexamples work even for the party-approval setting (Brill et al., 2022), where each candidate can be placed in the committee several times. For Phragmén, take the 3-voter profile $(ab,bc,ac)$, where Phragmén can elect $ababab$, with $T=\{c,c,c,c\}$ forming a deviation. For MES, take the 7-voter profile $(ab,ac,ad,bcd,bcd,bcd,bcd)$, where $bbbbbaa$ is an outcome of MES, with $T=\{c,c,c,d,d,d\}$ forming a deviation. (In these examples, there are other tied outcomes in the core, and I don’t know if unique examples exist.) Like other proofs that PAV is proportional, our proof reasons about the change in PAV score caused by certain swaps. To establish a key inequality, the proof refers to the result of a computer enumeration of all possible ballot types, which can be reproduced using Python code available on GitHub.
@@
When $k\leqslant 7$, every local PAV committee is in the core.
@@
We now consider the following process: we remove some candidate $x\in W\setminus T$ from the committee $W$, and replace it by some candidate $y\in T\setminus W$, and will estimate the total effect of these swaps. Because $W$ is a local PAV committee, we have
@@
$a=|(W\setminus T)\cap A|,\quad b=|(W\cap T)\cap A|,\quad c=|(T\setminus W)\cap A|.$
@@
While it may not look like it, the result of Theorem 4.1 was obtained by linear programming. Suppose that the result is false, so that there exists a profile $P$ and a local PAV committee $W$ such that some $T$ is a successful deviation. Without loss of generality, there exists such an example with $W=\{c_{1},\dots,c_{k}\}$. Note that $P$ then forms a solution to the following system of linear inequalities, where we may assume that $C=W\cup T$: [^5] We may make this assumption because a counterexample on a larger $C$ remains a counterexample when restricted to $W\cup T$, because a local PAV committee remains a local PAV committee after deleting candidates outside the committee.
@@
Due to the following remark, it is possible to compute a core-stable outcome in $O(m^{2}n)$ time whenever $k\leqslant 7$.
@@
Let $k\leqslant 7$ and $\varepsilon=0.1/k^{2}$. By solving linear programs [GitHub], one can check that every committee that is $\varepsilon$-local-swap-stable (i.e., $\Delta_{P,x,y}\leqslant\varepsilon$ for all $x$ and $y$) is in the core. An $\varepsilon$-local-swap-stable and thus a core-stable committee can be found by performing $O(k^{2}\ln k)=O(1)$ many $\varepsilon$-improving swaps (see Aziz et al., 2018, Proposition 1), each of which takes $O(m^{2}n)$ time to find. It turns out that for $\varepsilon=1/k^{2}$, it is not the case that every $\varepsilon$-local-swap-stable committee is in the core, even though this value of $\varepsilon$ is enough to ensure that the committee satisfies EJR (Aziz et al., 2018, Theorem 1). For example, in the profile with $P(\{a,b\})=P(\{a,c\})=0.25$ and $P(\{d,e,f,g,h\})=0.5$, for $k=6$, the committee $\{a,d,e,f,g,h\}$ fails the core due to $T=\{a,b,c\}$, but it is $1/40$-local-swap-stable, and $1/40<1/36=1/k^{2}$.
@@ 4.2 Committee size $k=8$
Theorem 4.1 does not hold for $k=8$: There are profiles where some local (and even global) PAV committee is not in the core.
@@
On this profile, $W=\{c_{1},c_{2},c_{5},c_{6},c_{7},c_{8},c_{9},c_{10}\}$ is a global PAV committee (indicated in blue in the picture). However, $W$ is not in the core: consider $T=\{c_{1},c_{2},c_{3},c_{4}\}$, which has support from $\frac{1}{2}$ of the voters, and $|T|/k=\frac{1}{2}$. ∎
Note, however, that in Example 4.3, there is more than one global PAV committee. In particular, $W^{\prime}=\{c_{1},c_{2},c_{3},c_{5},c_{6},c_{7},c_{8},c_{9}\}$ (obtained by removing $c_{10}$ and adding $c_{3}$) is also a global PAV committee and it is in the core.
@@
Let $P$ be a profile and suppose that $W$ with $|W|=8$ is a local PAV committee that is not in the core due to objection $T$. Then there exist distinct $a,b\in W$ and distinct $x,y\in C\setminus W$ such that $T=\{a,b,x,y\}$. In addition,
@@
We first check that if $W$ is not in the core, then any core objection must use a $T$ with $|T|=4$ and $|W\cap T|=2$. This can be deduced using the linear programming approach behind Theorem 4.1; by iterating through all possible $T$, we find that the system (4) has a solution only for $T$ satisfying the condition in the theorem statement. Alternatively, one can check [GitHub] that the inequality (3) is only violated for such $T$; thus for other $T$ the proof of Theorem 4.1 goes through.
Now fix such a $T=\{a,b,x,y\}$. Assume that there exists a profile $P$ where $W$ is a local PAV committee with successful deviation $T$ but that violates any of the conditions (i)–(iii). Then if we delete all candidates outside $W\cup T$ from the profile, it would still fail (i)–(iii). Thus, for purposes of making the following linear programs finite, we may assume that $C=W\cup T$ (so $|C|=10$).
@@
As we discussed, Example 4.3 shows an example of a global PAV committee that is not core-stable, but there are other global PAV committees for the same profile that are core-stable. Thanks to Lemma 4.4, we deduce that the same holds for all counterexamples. Hence, for every instance, at least one global PAV committee is in the core, and thus the core is always non-empty for $k=8$.
@@
When $k=8$, some global PAV committee is in the core.
@@
Let $k=8$ and let $P$ be a profile. If on $P$, all global PAV committees are in the core, we are done. So suppose that $W$ is a global PAV committee that is not core stable due to objection $T$. From Lemma 4.4, there exist distinct $a,b\in W$ and distinct $x,y\in C\setminus W$ such that $T=\{a,b,x,y\}$. Take any $c\in W\setminus\{a,b\}$. Then the committee $W^{\prime}=W\setminus\{c\}\cup\{x\}$ has the same PAV score as $W$, because the removal of $c$ causes a decrease in PAV score of $1/12$ and the addition of $x$ causes an increase of at least $\frac{1}{4}\cdot\frac{1}{3}=1/12$ due to the quarter of voters from (i) with ballots $A$ such that $A\cap(W\cup T)=\{a,b,x\}$. Thus, $W^{\prime}$ is also a global PAV committee. We now show that $W^{\prime}$ is in the core.